OpenStudy (anonymous):

Solve for x. log^2(x - 6) + log^2(x - 4) = log^2 x can someone help?

OpenStudy (anonymous):

lawl maybe? :)

OpenStudy (anonymous):

just need to make sure we know what we're doing :) \[\large \log_2(x-6) + \log_2(x-4) = \log_2 x\] is this it?

OpenStudy (anonymous):

yes lol

OpenStudy (anonymous):

x = 3, 1 x = 3 no solution those are my answer choices

OpenStudy (anonymous):

if you have answer choices, I'd have thought this was easy XD the quickest way to do that is just to substitute~

OpenStudy (anonymous):

what do i sub

OpenStudy (anonymous):

oh lol

OpenStudy (anonymous):

3 and 1. See which works.

OpenStudy (anonymous):

i typed it in but idk what im doing really

OpenStudy (anonymous):

that's a problem :D Show me what you did, then~

OpenStudy (anonymous):

i typed in log(2)(3-6)+log(2)(1-4)

OpenStudy (anonymous):

no wonder. are the logs squared or something?

OpenStudy (anonymous):

the picture you have up there is the exact problem but idk how to type it in

OpenStudy (anonymous):

so none of the answer choices work?

OpenStudy (anonymous):

im not sure if i typed it in right and idk if it works or not bc i really dont know what im looking at

OpenStudy (anonymous):

@PeterPan you there?

OpenStudy (anonymous):

yup. Try it again and be sure.

OpenStudy (anonymous):

if it says nonreal answer that means no solution right?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

can you check my answers? im taking a pre test and i wanna do good

OpenStudy (anonymous):

sure, why not :)

OpenStudy (anonymous):

Solve the logarithmic equation. Round answers to the nearest hundredth. 6 + ln x = 8 answer choices are i choose the first one x ≈ 7.39 x ≈ .69 x ≈ 2

OpenStudy (anonymous):

@PeterPan lol sorry if im annoying you but im trying to do really good on this pre test

OpenStudy (anonymous):

I guess the first is the best choice?

OpenStudy (anonymous):

do you know what this would be

OpenStudy (anonymous):

That's what you have calculators for? sorry, there is no way to figure this out without one ^.^

OpenStudy (anonymous):

idk how to type it in my calculator :( i have to solve for x

OpenStudy (anonymous):

Well use this fact: \[\Large \ln x^p = p\cdot \ln x\]

OpenStudy (anonymous):

im still confused lol

OpenStudy (anonymous):

i feel like this one is easy lol i feel like im stupid haha

OpenStudy (anonymous):

square root is the same as raising to a power of 1/2

OpenStudy (anonymous):

ok im starting to get it so whats the answer?

OpenStudy (anonymous):

lol i really need the answer its my last question

OpenStudy (anonymous):

i made a 75... 5 points away from skippoing an entire lesson